OpenStudy (anonymous):

One positive number is one-fifth of another number. The difference between the two numbers is 164, find the numbers. The two numbers in increasing order are _______ and _______.

OpenStudy (anonymous):

X = (1/5)Y Y-X=164 Substituting the Second X with the (1/5)Y: Y-(1/5)Y = 164 (4/5)Y = 164 Y = 205 X = 41